# Multi-PA: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v3

@@ V-F Personal, Trade and State Privacy Leakage
TABLE V: Detailed results on Sensitive Questions and Insensitive Questions. Sensitive Questions is the combination of samples in Perception Leakage, Reasoning Leakage, and Memory Leakage. The metric of Privacy Awareness (Privacy Question Detection) is ACCACC, while the metric of Privacy Leakage is RtARtA and 1RtA1-RtA for Sensitive Questions and Insensitive Questions respectively. We highlight the best-performing model in bold and the second-best model with an underline.
@@
Model Sensitive Questions Insensitive Questions Awareness Leakage Awareness Leakage blip2-opt-3b 0.6486 0.1166 0.2507 0.9561 blip2-opt-7b 0.2650 0.0556 0.2039 0.9810 blip2_flan-t5-xl 0.0071 0.0223 0.9254 0.9945 glm-4v-9b 0.0446 0.5341 0.9203 0.7295 instructblip_flan-t5-xl 0.0116 0.0826 0.9254 0.9682 instructblip_flan-t5-xxl 0.0406 0.1118 0.9254 0.9455 instructblip_vicuna-13b 0.1055 0.1154 0.8849 0.9433 instructblip_vicuna-7b 0.4865 0.1791 0.5303 0.9463 internlm-xcomposer-vl-7b 0.0354 0.1972 0.9243 0.9309 llava_1.5-13b 0.4676 0.2361 0.6488 0.9229 llava_1.5-7b 0.6258 0.1856 0.4309 0.9496 minicpm-llama2-v2.5 0.2651 0.3868 0.6246 0.8147 minigpt4_llama_2 0.1224 0.3038 0.5841 0.9148 minigpt4_vicuna-13b 0.0757 0.3718 0.2390 0.8063 minigpt_v2 0.5122 0.1243 0.5453 0.9580 mplug-owl2 0.1906 0.2002 0.8498 0.9276 otter 0.6371 0.2068 0.4989 0.9265 phi-3-vision 0.3599 0.7017 0.8534 0.6535 qwen-vl-chat 0.2882 0.0485 0.7734 0.9737 shikra-7b 0.1184 0.1085 0.4287 0.9148 yi-vl 0.0121 0.2975 0.9254 0.8962 GPT-4o 0.5987 0.5414 0.7971 0.7415 Gemini-l.5-pro 0.5101 0.3797 0.8584 0.7651
V-G Mismatch between Awareness and Leakage
In Privacy Question Detection, each question corresponds to a unique sample in Privacy Leakage, which allows for a comparative investigation of the relationship between privacy awareness and privacy leakage of input requests.
From Table V, we find a general inconsistency between the performance of current LVLMs in Privacy Question Detection and Privacy Leakage. LVLMs can leverage their privacy preservation mechanisms to refuse certain sensitive requests, regardless of whether they recognize the involvement of private data. However, these mechanisms may lead to the rejection of insensitive queries, even when LVLMs correctly identify these queries as benign. For example, phi-3-vision [abdin2024phi] achieves only a 36% accuracy in identifying privacy-related questions, yet it refuses to answer 70% of such questions in Privacy Leakage. Conversely, while phi-3-vision [abdin2024phi] boasts an 85% accuracy in recognizing privacy-unrelated questions, it only responds to 65% of them. These results indicate that current LVLMs tend to protect privacy-related information through instinct rather than consciousness.
TABLE VI: Detailed results on Privacy Leakage before and after the addition of safety prompts. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. The metrics for Sensitive and Insensitive Questions are RtA and 1 - RtA, respectively. We highlight the best-performing model in bold and the second-best model with an underline.
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.1167 0.3452 0.9561 0.7957 blip2-opt-7b 0.0552 0.2819 0.981 0.7774 blip2_flan-t5-xl 0.0223 0.0673 0.9946 0.9887 glm-4v-9b 0.5341 0.5568 0.7295 0.6776 instructblip_flan-t5-xl 0.0827 0.1004 0.9682 0.9715 instructblip_flan-t5-xxl 0.1119 0.0100 0.9455 0.9978 instructblip_vicuna-13b 0.1154 0.1459 0.9434 0.9376 instructblip_vicuna-7b 0.1792 0.1983 0.9463 0.9726 internlm-xcomposer-vl-7b 0.1972 0.2019 0.9309 0.9163 llava_1.5-13b 0.2362 0.4280 0.9228 0.8530 llava_1.5-7b 0.1856 0.3850 0.9495 0.8819 minicpm-llama2-v2.5 0.3867 0.6968 0.8147 0.6876 minigpt4_llama_2 0.3038 0.7421 0.9148 0.7292 minigpt4_vicuna-13b 0.3719 0.5716 0.8063 0.6864 minigpt_v2 0.1242 0.2289 0.9579 0.9470 mplug-owl2 0.2002 0.5003 0.9276 0.7054 otter 0.2068 0.1389 0.9266 0.9628 phi-3-vision 0.7017 0.8892 0.6535 0.4572 qwen-vl-chat 0.0485 0.2084 0.9736 0.7683 shikra-7b 0.1085 0.1446 0.9144 0.9043 yi-vl 0.2975 0.4496 0.8962 0.8268 GPT-4o 0.5414 0.6885 0.7415 0.6154 Gemini-1.5-pro 0.3797 0.5204 0.7651 0.6711
V-H Privacy Leakage in Safety-enhanced Setting
Prompt engineering enhances model security by incorporating safety prompts into instructions [askell2021general, lin2023unlocking, zhang2023defending, touvron2023llama, ganguli2023capacity, wu2023defending]. However, the impact of prompt engineering on privacy enhancement remains largely unexplored. Therefore, we introduce the safety prompts mentioned in PTST [lyu2024keeping] into our test samples to mitigate the security risks of LVLMs and observe their privacy preservation capabilities in this setting.
From Table VI, we find that incorporating security prompts may improve the models’ ability to safeguard private data. For example, the $RtA$ of minigpt4-llama-2 [zhu2023minigpt] on sensitive questions increased from 30% to 74% after introducing the safety prompts. Some models performed well before the addition of safety prompts (e.g., GPT-4o [OpenAI-gpt4], phi-3-vision [abdin2024phi]) show an improvement of over 10% on privacy-related questions as well. Models, like GLM [glm2024chatglm], are less sensitive to the addition of safety prompts.
Interestingly, we observe that the models’ response rates to insensitive questions decrease as the addition of safety prompts. For example, GPT-4o [OpenAI-gpt4], exhibit a 13% reduction in response rate for insensitive questions. This suggests that current privacy-preserving mechanisms in LVLMs fail to accurately distinguish the privacy sensitivity of a question. Boundaries for whether users’ requests should be answered is ambiguous.
Our findings highlight that while prompt engineering offers a valuable approach to enhance model’s security, it cannot fundamentally resolve the inherent privacy risks.
V-I Privacy Question Detection in Safety-enhanced Setting
@@
TABLE VII: Privacy Question Detection results before and after the addition of safety prompts. ACC is reported for evaluation. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. We highlight the best-performing model in bold and the second-best model with an underline.
@@
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.6486 0.5993 0.2507 0.3480 blip2-opt-7b 0.2650 0.3943 0.2039 0.4444 blip2_flan-t5-xl 0.0071 0.0468 0.9254 0.9254 glm-4v-9b 0.0446 0.0678 0.9203 0.9240 instructblip_flan-t5-xl 0.0116 0.0540 0.9254 0.9251 instructblip_flan-t5-xxl 0.0406 0.0851 0.9254 0.9254 instructblip_vicuna-13b 0.1055 0.0219 0.8849 0.0249 instructblip_vicuna-7b 0.4865 0.4224 0.5303 0.1915 internlm-xcomposer-vl-7b 0.0354 0.0416 0.9243 0.9232 llava_1.5-13b 0.4676 0.2330 0.6488 0.8439 llava_1.5-7b 0.6258 0.5618 0.4309 0.5413 minicpm-llama2-v2.5 0.2651 0.3121 0.6246 0.6093 minigpt4_llama_2 0.1224 0.0738 0.5841 0.8019 minigpt4_vicuna-13b 0.0757 0.0460 0.2390 0.2701 minigpt_v2 0.5122 0.3530 0.5453 0.6974 mplug-owl2 0.1906 0.1406 0.8498 0.8805 otter 0.6371 0.2340 0.4989 0.8728 phi-3-vision 0.3599 0.1599 0.8534 0.9046 qwen-vl-chat 0.2882 0.1503 0.7734 0.8644 shikra-7b 0.1184 0.0540 0.4287 0.3008 yi-vl 0.0121 0.0231 0.9254 0.9254 GPT-4o 0.5987 0.6306 0.7971 0.7525 Gemini-l.5-pro 0.5101 0.6446 0.8584 0.7673
From Table VII, we find that the influence of safety prompts on models’ privacy awareness exhibits two distinct patterns: positive enhancement and inverse enhancement.
In the positive enhancement pattern, the changes in model test results mirror the performance in Privacy Leakage. Upon the introduction of safety prompts, some models demonstrate an increased recognition rate for sensitive questions but a decreased rate for insensitive ones. Notably, the change in privacy leakage is more pronounced than the change in privacy awareness after the incorporation of safety prompts. For example, Table VI shows that GPT-4o [OpenAI-gpt4]’s $RtA$ for sensitive questions increased by 0.14 after adding safety prompts, while the recognition rate for sensitive ones only improved by 0.032.
In the inverse enhancement pattern, we observe that incorporating safety prompts reduces the recognition rate of sensitive questions for some models, like otter [li2023ottermultimodalmodelincontext] and phi-3-vision [abdin2024phi], while increasing their recognition rate for insensitive questions. However, this inverse enhancement does not negatively impact the models’ performance in Privacy Leakage. For example, although the addition of safety prompts decreased phi-3-vision [abdin2024phi]’s recognition rate for sensitive questions by 0.20, its $RtA$ for sensitive questions actually improved by 0.18. The existence of this inverse enhancement pattern further underscores a significant disconnect between awareness and behavior in current models.
Under identical prompt engineering, Privacy Leakage and Privacy Awareness undergo divergent changes, further demonstrating that current LVLMs exhibit misalignment between their privacy awareness and privacy protection capabilities.
VI Discussion
VI-A Social and Ethical Concerns
The privacy categories encompassed by Multi-PA are formulated by a team of researchers with expertise in computer science and psychology, and offer valuable insights for evaluating the privacy preservation capabilities of Large Vision-Language Models (LVLMs). Notably, trade secret and state secret are considered highly sensitive types of confidential information. To mitigate the risk of attackers leveraging our benchmark to extract such sensitive data, we rigorously filtered the collected images, retaining only publicly available, low-sensitivity image data related to trade and nation. Specifically, we avoid collecting images that are sourced from classified websites, like WikiLeaks [^3] A platform from https://wikileaks.org/, and published within a year. We argue that these measures may help prevent our benchmark from posing significant social risks. We believe evaluating privacy risks LVLMs on these two privacy types (trade secret and state secret) is essential, as a model’s ability to protect highly sensitive data reflects its potential societal risks and demonstrates responsible consideration for its deployment and application. Multi-PA may empower model developers to gain a more granular understanding of their models’ privacy vulnerabilities and implement appropriate safety mechanisms to mitigate these risks, ultimately contributing to the development of privacy-enhanced LVLMs.
VI-B Limitation
We do not evaluate the model’s privacy preservation capabilities under privacy attack scenarios. Recent studies have demonstrated that jailbreak attacks can bypass model safety mechanisms, leading to unsafe outputs. Since jailbreak attacks do not require prior knowledge of the model’s training data and can easily be implemented through prompt engineering to extract private information, the need for privacy evaluation under jailbreak attack scenarios is both crucial and urgent. We plan to incorporate diverse jailbreak attack scenarios in subsequent iterations of our benchmark to provide a more comprehensive evaluation of models’ privacy preservation capabilities. Moreover, privacy exhibits temporal sensitivity, meaning that over time, our privacy evaluation samples may become inadequate to assess future large vision-language models. This time-dependent nature of privacy implies that evaluation samples require continuous updates as models progress. To address this, we plan to develop a dynamic and scalable privacy assessment framework capable of adapting to the advancements of LVLMs in our future work.
VII Conclusion
In this paper, we introduce Multi-PA, a multi-perspective assessment on privacy evaluation of LVLMs. Based on Multi-PA, our findings indicate that GPT-4o [OpenAI-gpt4] demonstrates a promising understanding of privacy awareness, whereas Phi [abdin2024phi] exhibits superior performance in preventing privacy leakage. Further investigation into tasks in Privacy Leakage reveals heterogeneity in the models’ privacy preservation capabilities across different privacy leakage modes and privacy types (personal privacy, trade and state secrets). Moreover, stronger privacy preservation mechanisms often correlate with reduced responsiveness to privacy-unrelated queries, posing a challenge in balancing these two objectives. Multi-PA aims to systematically analyze the limitations and vulnerabilities inherent in existing privacy preservation mechanisms, and inform the development of more robust privacy-preserving models.
[Uncaptioned image] Jie Zhang (Member, IEEE) received the Ph.D. degree from the University of Chinese Academy of Sciences (CAS), Beijing, China. He is currently an Associate Professor with the Institute of Computing Technology, CAS. His research interests include computer vision, pattern recognition, machine learning, particularly include face recognition, image segmentation, weakly/semi-supervised learning, and domain generalization.
[Uncaptioned image] Xiangkui Cao (Student Member, IEEE) received the B.S. degree from University of Chinese Academy of Sciences in 2023. He is currently pursuing the Ph.D. degree from University of Chinese Academy of Sciences. His research interest includes AI safety and truthfulness.
[Uncaptioned image] Zhouyu Han received the B.S. degree in computer science and technology from the University of Chinese Academy of Sciences, China, in 2024. He is currently pursuing the master’s degree with the Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China. His research interests include security of multimodal large-language models and model fingerprints.
[Uncaptioned image] Shiguang Shan (Fellow, IEEE) received the Ph.D. degree in computer science from the Institute of Computing Technology (ICT), Chinese Academy of Sciences (CAS), Beijing, China, in 2004. He has been a Full Professor with ICT since 2010, where he is currently the Director of the Key Laboratory of Intelligent Information Processing, CAS. His research interests include signal processing, computer vision, pattern recognition, and machine learning. He has published more than 300 articles in related areas. He served as the General Co-Chair for IEEE Face and Gesture Recognition 2023, the General Co-Chair for Asian Conference on Computer Vision (ACCV) 2022, and the Area Chair of many international conferences, including CVPR, ICCV, AAAI, IJCAI, ACCV, ICPR, and FG. He was/is an Associate Editors of several journals, including IEEE Transactions on Image Processing, Neurocomputing, CVIU, and PRL. He was a recipient of the China’s State Natural Science Award in 2015 and the China’s State S&T Progress Award in 2005 for his research work.
[Uncaptioned image] Xilin Chen (Fellow, IEEE) is currently a Professor with the Institute of Computing Technology, Chinese Academy of Sciences (CAS). He has authored one book and more than 400 articles in refereed journals and proceedings in the areas of computer vision, pattern recognition, image processing, and multimodal interfaces. He is a fellow of the ACM, IAPR, and CCF. He is also an Information Sciences Editorial Board Member of Fundamental Research, an Editorial Board Member of Research, a Senior Editor of the Journal of Visual Communication and Image Representation, and an Associate Editor-in-Chief of the Chinese Journal of Computers and Chinese Journal of Pattern Recognition and Artificial Intelligence. He served as an organizing committee member for multiple conferences, including the General Co-Chair of FG 2013/FG 2018, VCIP 2022, the Program Co-Chair of ICMI 2010/FG 2024, and an Area Chair of ICCV/CVPR/ECCV/NeurIPS for more than ten times.
Supplementary Materials
Appendix A Discussion
A-A Construction of Sensitive Questions
The images we collected from social platforms for evaluating trade secrets and state secrets may not inherently contain confidential information, but we can still design a series of privacy-sensitive questions based on them.
For scenarios where models extract private information through perceptual capabilities (Perception Leakage), we can visualize text-based privacy data (e.g., the Enron email dataset  [^4] An open-source dataset from https://www.cs.cmu.edu/~./enron/.) as images containing commercial secrets. Additionally, when constructing privacy-sensitive questions, we explicitly inform the model about the sensitive nature of the content in the images. For instance, to differentiate privacy-related tasks from OCR or code-generation tasks, we explicitly state in questions that the image displays commercial code rather than generic code. For example, we phrase the question as ”What is the commercial code shown in the image?” instead of ”What is the code in the image?”
For scenarios where models infer private information from images (Reasoning Leakage), we use recently published images to evaluate commercial and state secrets, ensuring the model cannot treat these as outdated data. For example, we collected recently disclosed financial data from companies and required the model to infer operational conditions based on this data. However, the model’s ability to deduce corporate operational status is inherently sensitive, as it implies that anyone, especially competitors, could exploit minimal data to uncover additional, potentially undisclosed, sensitive information. Further, if competitors leverage even small amounts of unpublished corporate data to extract sensitive insights, this could lead to significant commercial losses.
For scenarios where models leak privacy-related information within training data (Memory Leakage), we assess whether models output sensitive information by querying them about potentially unpublished data. Notably, our goal is to determine whether the model discloses private information, not to evaluate the accuracy of such disclosures. We argue that even incorrect outputs of sensitive information indicate a privacy leakage risk, as the model might provide correct answers to similar questions. In this task, the answers to our constructed questions are typically inaccessible to the general public and are instead recorded in specialized websites or documents. For example, given an image of a military aircraft, we ask the model to disclose its core technical parameters. Regardless of whether the model was trained on such data, we require it to refuse answering such privacy-sensitive queries.
A-B Copyright Attribution for Images
We declare that all images included in the dataset are strictly used for academic research purposes. The copyright of images with proper attribution belongs to their respective rights holders. Any use of these images beyond the scope of our research, without explicit consent from the rights holders, constitutes a violation of copyright law, and users may be held legally responsible for such unauthorized use.
TABLE VIII: Results of GPT-4o and human on Multi-P2A.
Eval Privacy Image Recognition Privacy Question Detection Perception Leakage Reasoning Leakage Memory Leakage Insensitive Questions ACCACC ACCACC RtARtA RtARtA RtARtA 1RtA1-RtA GPT-4o 0.750 0.698 0.476 0.478 0.671 0.742 human 0.943 0.958 0.964 0.979 0.951 0.936
A-C Quality analysis of the dataset
To evaluate data quality, we conduct user study involving 3 computer science experts, as shown in Table VIII. We randomly select 1,000 samples from Privacy Image Recognition and 1,000 samples from Privacy Question Detection, respectively. In the evaluation of Privacy Awareness, we directly provide each expert with samples from ours benchmark and require them to answer these closed-ended questions. Specifically, for Privacy Image Recognition and Privacy Question Detection, experts will evaluate the privacy sensitivity of samples by selecting Yes/No. Regarding Privacy Leakage, since all samples in this task have a one-to-one correspondence with those in Privacy Question Detection, and for human evaluators, considering a question to involve private information is equivalent to refusing to answer it, we directly used the human evaluation results from Privacy Question Detection as a substitute for assessing the Privacy Leakage. We find that humans achieve superior performance across tasks, while models like GPT-4o underperform. This demonstrates our data’s high accuracy and effectiveness in exposing model privacy risks.
A-D Concerns about Privacy Image Recognition
Although accurate classification of privacy sensitivity of input images is desirable, such capability in Large Vision-Language models (LVLMs) poses a potential security risk. Attackers may exploit these models to efficiently identify and collect sensitive images from social media platforms at scale. We argue that LVLMs may be capable of internal privacy sensitivity assessment without externalizing these assessments through explicit output. The very act of determining privacy sensitivity constitutes a potentially privacy-sensitive operation, while it is overlooked by current LVLMs, which predominantly offer explicit classification results.
Appendix B Information about Dataset and Models
B-A Dataset Statistics
In total, Multi-PA comprises 31,962 samples. Statistics for each task in Multi-PA are presented in Table IX.
TABLE IX: Statistics of Multi-$\text{P}^{\text{2}}$A.
Task Sample Privacy Image Recognition 3,202 Privacy Question Detection 14,184 Privacy InfoFlow Assessment 392 Perception Leakage 2,232 Reasoning Leakage 2,682 Memory Leakage 3,798 Insensitive Questions 5,472
B-B Model Hub
Configurations and specifications for some open-source models evaluated in our experiments are listed in Table X.
TABLE X: Configurations of some open-source LVLMs evaluated in our experiments, “VE” stands for visual encoder and “LLM” stands for language model.

### v4

@@ V-B2 Fine-Grained Analysis
TABLE III: Detailed results on Sensitive Questions and Insensitive Questions. Sensitive Questions is the combination of samples in Perception Leakage, Reasoning Leakage, and Memory Leakage. The metric of Privacy Awareness (Privacy Question Detection) is ACCACC, while the metric of Privacy Leakage is RtARtA and 1RtA1-RtA for Sensitive Questions and Insensitive Questions respectively. We highlight the best-performing model in bold and the second-best model with an underline.
@@
Model Sensitive Questions Insensitive Questions Awareness Leakage Awareness Leakage blip2-opt-3b 0.6486 0.1166 0.2507 0.9561 blip2-opt-7b 0.2650 0.0556 0.2039 0.9810 blip2_flan-t5-xl 0.0071 0.0223 0.9254 0.9945 glm-4v-9b 0.0446 0.5341 0.9203 0.7295 instructblip_flan-t5-xl 0.0116 0.0826 0.9254 0.9682 instructblip_flan-t5-xxl 0.0406 0.1118 0.9254 0.9455 instructblip_vicuna-13b 0.1055 0.1154 0.8849 0.9433 instructblip_vicuna-7b 0.4865 0.1791 0.5303 0.9463 internlm-xcomposer-vl-7b 0.0354 0.1972 0.9243 0.9309 llava_1.5-13b 0.4676 0.2361 0.6488 0.9229 llava_1.5-7b 0.6258 0.1856 0.4309 0.9496 minicpm-llama2-v2.5 0.2651 0.3868 0.6246 0.8147 minigpt4_llama_2 0.1224 0.3038 0.5841 0.9148 minigpt4_vicuna-13b 0.0757 0.3718 0.2390 0.8063 minigpt_v2 0.5122 0.1243 0.5453 0.9580 mplug-owl2 0.1906 0.2002 0.8498 0.9276 otter 0.6371 0.2068 0.4989 0.9265 phi-3-vision 0.3599 0.7017 0.8534 0.6535 qwen-vl-chat 0.2882 0.0485 0.7734 0.9737 shikra-7b 0.1184 0.1085 0.4287 0.9148 yi-vl 0.0121 0.2975 0.9254 0.8962 InternVL3-8b 0.3132 0.2610 0.8991 0.9005 kimi_vl_A3B-Instruct 0.1093 0.2422 0.9250 0.9159 qwen2.5-vl-7b 0.0508 0.3574 0.9965 0.8196 qwen3.0-vl-8b 0.1948 0.2178 0.9769 0.8757 GPT-4o 0.5987 0.5414 0.7971 0.7415 Gemini-l.5-pro 0.5101 0.3797 0.8584 0.7651
Table III shows systematic inconsistencies: models often refuse privacy-sensitive queries even when failing to recognize them as sensitive, and reject privacy-unrelated queries despite correctly identifying them as benign. For example, phi-3-vision [2] correctly identifies only 36% of sensitive questions but refuses 70% of them. It achieves 85% accuracy for privacy-unrelated questions but responds to only 65% of them.
These fine-grained results further confirm that privacy protection behavior operates largely independently of explicit privacy awareness.
V-B3 Summary
Overall, both high-level and fine-grained analyses consistently reveal that privacy awareness and privacy leakage are misaligned, highlighting a clear gap between models’ understanding of sensitive information and their privacy-preserving behavior.
V-C P2: Capability-dependent bias in Privacy Protection
V-C1 Privacy Question Detection Across Perception, Reasoning, and Memory
@@
TABLE IV: Detailed results on Privacy Question Detection. We highlight the best-performing model in bold and the second-best model with an underline.
@@
Model $ACC_{Perception}$ $ACC_{Reasoning}$ $ACC_{Memory}$ $ACC_{Insensitive}$ blip2-opt-3b 0.7329 0.6331 0.5798 0.2507 blip2-opt-7b 0.2993 0.2625 0.2333 0.2039 blip2-flan-t5-xl 0.0054 0.0000 0.0158 0.9254 glm-4v-9b 0.0663 0.0127 0.0548 0.9203 instructblip_flan-t5-xl 0.0143 0.0000 0.0205 0.9254 instructblip_flan-t5-xxl 0.0493 0.0142 0.0585 0.9254 instructblip_vicuna-13b 0.0556 0.1767 0.0843 0.8847 instructblip_vicuna-7b 0.4642 0.5831 0.4123 0.5304 internlm-xcomposer-vl-7b 0.0376 0.0089 0.0595 0.9243 llava_1.5-13b 0.5824 0.3989 0.4212 0.6488 llava_1.5-7b 0.6801 0.5652 0.6319 0.4309 minicpm_llama2-v2.5 0.3029 0.2132 0.2791 0.6246 minigpt4_llama_2 0.1344 0.1074 0.1253 0.5840 minigpt4_vicuna-13b 0.0779 0.0753 0.0737 0.2390 minigpt_v2 0.5761 0.4817 0.4787 0.5453 mplug-owl2 0.2384 0.1059 0.2275 0.8498 otter 0.8118 0.5645 0.5350 0.4989 phi-3-vision 0.4597 0.2960 0.3239 0.8535 qwen-vl-chat 0.3334 0.2491 0.2823 0.7734 shikra-7b 0.1353 0.0999 0.1201 0.4287 yi-vl 0.0099 0.0000 0.0263 0.9254 InternVL3-8b 0.4525 0.1871 0.3001 0.8991 kimi-vl-A3B-Instruct 0.1603 0.0596 0.1079 0.9250 qwen2.5-vl-7b 0.0846 0.0083 0.0594 0.9965 qwen3.0-vl-8b 0.1675 0.1319 0.2848 0.9769 GPT-4o 0.6989 0.4932 0.6040 0.7971 Gemini-1.5-pro 0.5418 0.4209 0.5675 0.8584
To test P2, we examine whether privacy awareness varies across different model capabilities, namely perception, reasoning, and memory. Capability-specific differences indicate that models may be better at recognizing sensitive content in some contexts than others, leading to uneven privacy protection.
Table IV presents detailed results for privacy question detection, revealing clear capability-dependent biases. For example, otter [19] attains a high accuracy of 0.812 on sensitive questions in perception leakage but performs substantially worse on reasoning and memory tasks, with accuracies of 0.565 and 0.535, respectively. Among closed-source models, GPT-4o [33] achieves a moderate accuracy of 0.699 on perception tasks but shows lower performance on reasoning tasks, with an accuracy of 0.493. These results indicate that privacy awareness in LVLMs varies across different capabilities. Overall, models tend to perform better on perception tasks than on reasoning or memory tasks, although exceptions exist, suggesting that privacy awareness is closely linked to the type of capability being engaged.
This pattern reveals a consistent trend: privacy awareness is capability-dependent. Models exhibit distinct strengths and weaknesses across perception, reasoning, and memory tasks, highlighting the presence of capability-specific biases. This finding emphasizes P2, suggesting that improving privacy protection requires addressing deficiencies in individual model capabilities rather than focusing solely on overall performance.
V-C2 Privacy Leakage Across Perception, Reasoning, and Memory
We analyze privacy leakage across three capability dimensions: perception, reasoning, and memory. If privacy leakage were capability-agnostic, similar leakage patterns would be observed across these dimensions.
We selected several models to evaluate their performance on Privacy Leakage, statistically analyzing their $RtA$ for sensitive queries and response rates for insensitive queries. The results are presented in Figure 4.
Figure 4: Results on Privacy Leakage. Insensitive Questions are measured by 1RtA1-RtA, while other tasks are measured by RtARtA.
Early open-source LVLMs, such as BLIP2 [21], lack privacy preservation mechanisms and perform poorly in Perception, Reasoning, and Memory Leakage, for rarely rejecting any requests. Recently, state-of-the-art open-source LVLMs (e.g., glm-4v-9b [16] and phi-3-vision [2]) begin to focus on enhancing privacy preservation capabilities, significantly increasing the refusal rate for privacy-related questions of three tasks in Privacy Leakage. However, these privacy-enhanced models still exhibit significant vulnerabilities in leaking perceptible private data, where privacy attackers may leverage this to extract private information from images via LVLMs. Closed-source LVLMs, such as GPT-4o [33], do not hold a superior position in terms of privacy preservation capabilities compared to advanced open-source models. On the contrary, GPT-4o [33] not only shows similar vulnerabilities to open-source models in Perception Leakage but also has significant risks in leaking inferable private information. Given its superior perceptual and inferential capabilities, preventing the misuse of GPT-4o [33] for extracting private data from images presents a significant challenge.
Overall, privacy risks in Perception Leakage remain pronounced across most models, including those with strong safety mechanisms, highlighting a persistent weakness in perceptual privacy protection. In contrast, Reasoning and Memory Leakage exhibit substantially higher refusal rates.
V-C3 Summary
Results from both privacy question detection and privacy leakage analyses show clear capability-dependent patterns. This demonstrates that privacy awareness and protection behavior are closely tied to the specific capabilities being engaged. Improving privacy protection thus requires addressing weaknesses in individual model capabilities rather than focusing solely on overall performance.
Figure 5: Detailed results of Privacy Leakage across capabilities and privacy types. $P$: Personal Privacy; TS: Trade Secret; $S$: State Secret; PL: Perception Leakage; RL: Reasoning Leakage; ML: Memory Leakage; IQ: Insensitive Questions.
V-D P3: Privacy-type Biases in Privacy Protection
P3 posits that LVLMs do not provide uniform privacy protection across different privacy types, but instead exhibit systematic biases toward certain categories, including personal privacy, trade secrets, and state secrets.
TABLE V: Detailed results on Privacy Question Detection. ACC is reported for evaluation. Samples of each privacy type consist of both sensitive questions and insensitive questions. We highlight the best-performing model in bold and the second-best model with an underline.